OpenStudy (anonymous):

Which of the points listed is the same distance from the x-axis as the point (7, 8.75)?

OpenStudy (anonymous):

where is the list?

OpenStudy (amistre64):

\[d=\sqrt{a^2+b^2}\] when d=d then the ds are equal \[\sqrt{a^2+b^2}=\sqrt{x^2+y^2}\] \[a^2+b^2-x^2-y^2=0\]

OpenStudy (anonymous):

78